Ms. Ellis, a single individual, had $115,000 taxable income. Assume the taxable year is 2018. Compute her income tax assuming that:

A) Taxable income includes no capital gain.

B) Taxable income includes $22,000 capital gain eligible for the 15 percent preferential rate
Answer #1

Solution:

Computation of income tax of Ms. Ellis

a) When taxable on come does not include capital gain and taxable income is $115,000

If the taxable income is $115,000 then Ms. Ellis will fall under the tax bracket of

Over $91,150 but not over $190,150

Tax to be paid = $18,558.75 +28% in excess of $91,150

= 18,558.75 + 28% (115,000 - 91,150)

=188,558.75 + 6,678

Tax to be paid by Ms. Ellis = $ 25,236.75

b)when taxable income does include capital gain of 22,000 eligible for preference rate of 15% and taxable income id $115,000

If the taxable income includes capital gain , then taxable income without capital gain is $93,000 ($115,000 - $22,000)

If the taxable income is $93,000 then Ms. Ellis will fall under the tax bracket of over $91,150 but not over $190,150

Tax to be paid = $18,558.75 + 28% in excess of $91,150

= 18,558.75 + 28% (93,000 - 91,150)

= 18,558.75 + 518

tax on total income other than capital gain = $19,076.75

Tax on income from capital gain = income from capital gain * tax rate = 22,000 * 15%

Tax on income from capital gain = $3,300

Therefore total tax to be paid = $19,076.75 + $3,300

Tax to be paid by Ms Ellis = $ 22,376.75